### v2

@@ 2 The housing market model
An allocation is a one-to-one mapping $\mu$ from $I$ to $O$ where $\mu(i)$ denotes the object received by $i$. An allocation $\mu$ is individually rational if, $\forall i\in I$, $\mu(i){}_{i}\omega(i)$. Two allocations $\mu$ and $\mu^{\prime}$ are (welfare) equivalent if, $\forall i\in I$, $\mu(i)\sim_{i}\mu^{\prime}(i)$. An allocation $\sigma$ is a Pareto improvement over $\mu$ for $C$ if, $\forall i\in C$, $\sigma(i){}_{i}\mu(i)$, and for some $j\in C$, $\sigma(j)\succ_{j}\mu(j)$. An allocation is Pareto efficient if it does not have a Pareto improvement for all agents.
We define some useful notations. Given any $i\in I$, for every $O^{\prime}\subseteq O$, let $B_{i}(O^{\prime})=\{o\in O^{\prime}:o{}_{i}o^{\prime}\text{ for all }o^{\prime}\in O^{\prime}\}$, the set of $i$’s favorite objects among $O^{\prime}$. For every $o\in O$, let $\mathcal{I}_{i}(o)=\{o^{\prime}\in O:o^{\prime}\sim_{i}o\}$, the set of objects $i$ views as indifferent with $o$. For any coalition $C$ and any allocation $\mu$, let $\mu(C)=\cup_{i\in C}\{\mu(i)\}$, the set of objects assigned to $C$. Then, $\omega(C)$ is the set of objects owned by $C$. For any two allocations $\mu$ and $\sigma$, we call an agent unaffected from $\mu$ to $\sigma$ if he receives indifferent objects in the two allocations. Then, we let $C_{\sigma>\mu}=\{i\in C:\sigma(i)\succ_{i}\mu(i)\}$ and $C_{\sigma\sim\mu}=\{i\in C:\sigma(i)\sim_{i}\mu(i)\}$, respectively the set of members of $C$ who become strictly better off from $\mu$ to $\sigma$ and the set of members of $C$ whose welfare remains unaffected.
[…]
We introduce the following property as a criterion to assess solution concepts. It requires that a solution is “closed” regarding equivalent allocations: if one allocation belongs to the solution, any other equivalent allocation also belongs to the solution. This property appears intuitive for coalition blocking, as agents should base their blocking decisions solely on their welfare.
@@
[…]
3 Existing solutions: standard core and exclusion core
@@ 3.1 Strong core and weak core
In the standard definition, a coalition blocks an allocation if its members can benefit from a reallocation of their endowments among themselves. Depending on whether all members must be strictly better off in the deviation, the definition has two variants.
@@
In the above definition, if $\forall i\in C$, $\sigma(i)\succ_{i}\mu(i)$, then we say that $\mu$ is strongly blocked by $C$ via $\sigma$. The weak core consists of allocations that are not strongly blocked.
@@
The strong core is a subset of the weak core. Under strict preferences, the strong core is nonempty and a singleton. Under weak preferences, however, the strong core may be empty, but if nonempty, it must be Pareto efficient. No matter preferences are strict or not, the weak core is nonempty but may not be Pareto efficient. This difference arises because unaffected agents are excluded from strong blocking coalitions, yet their participation is necessary for ensuring Pareto efficiency. See the following example for an illustration.
@@
Three agents $1,2,3$ respectively own three objects $a,b,c$. In the following tables, the left one lists agents’ endowments and the allocations under our examination. The other two tables represent two preference profiles.
@@
(Strict preferences) In I, since agents’ favorite objects are distinct, $\sigma$ is the unique Pareto efficient allocation. The weak core = $\{\mu,\sigma\}$, while the strong core = $\{\sigma\}$. $\mu$ is weakly blocked by $\{1,2,3\}$ via $\sigma$, and $2$ is unaffected from $\mu$ to $\sigma$. In contrast, $\mu$ is not strongly blocked because, given that $2$ cannot be made strictly better off and thus cannot join any blocking coalition, $\{1,3\}$ have no rights to reallocate $b$.
(Weak preferences) In $^{\prime}_{I}$, both $1$ and $3$ most prefer $2$’s endowment and then prefer their own endowments, while $2$ is indifferent between all objects. Thus, $\mu$ and $\sigma^{\prime}$ are the individually rational and Pareto efficient allocations for $^{\prime}_{I}$. In $\mu$, $1$ and $2$ exchange endowments, while in $\sigma^{\prime}$, $3$ and $2$ exchange endowments.
The weak core = $\{\omega,\mu,\sigma^{\prime}\}$. $\omega$ is not Pareto efficient, but it is not strongly blocked because $2$ cannot be made strictly better off and thus cannot join any blocking coalition.
In contrast, the strong core = $\emptyset$. $\mu$ and $\sigma^{\prime}$ are the only candidates for the strong core. But $\mu$ is weakly blocked by $\{2,3\}$ via $\sigma^{\prime}$, while $\sigma^{\prime}$ is weakly blocked by $\{1,2\}$ via $\mu$. Because $2$ is indifferent between $a$ and $c$, 22 alternates between $1$ and $3$ in forming weak blocking coalitions.
Under weak preferences, quint2004houseswapping provide a necessary and sufficient condition for the strong core to be nonempty. Their condition uses the following algorithm to partition agents into groups.
Partition by minimal self-mapped sets (PMSS)
@@
A minimal self-mapped set is a group of agents for whom all of their most preferred objects are owned by group, and no strict subset of the group satisfies the same condition. Given a market, let $t^{*}$ denote the last step of PMSS and $T^{*}=(T_{1},T_{2},\ldots,T_{t^{*}})$ denote the sequence of minimal self-mapped sets generated in PMSS.
Definition 3 (quint2004houseswapping).
@@
In words, a TTS is a sequence of minimal self-mapped sets $(T_{1},T_{2},\ldots,T_{t^{*}})$ in which each $T_{k}$ can distribute their objects among themselves in a way such that each $i\in T_{k}$ obtains one of his most preferred objects among those owned by $T_{k}\cup T_{k+1}\cup\cdots\cup T_{t^{*}}$.
Proposition 0 (quint2004houseswapping).
[…]
A corollary of this result is that, when the strong core is nonempty, all of its elements are equivalent, and any allocation equivalent to the elements of the strong core belongs to the strong core. This implies that the strong core is equivalence-closed. It is easy to prove that the weak core is also equivalence-closed (see Proposition 2).
3.2 Balbuzanov and Kotowski’s exclusion core
In a general model encompassing the housing market model as a special case, assuming strict preferences, balbuzanov2019endowments propose the exclusion core as a new solution concept. The exclusion core and the strong core do not include each other in the general model, but they coincide in the housing market model under strict preferences. When preferences are weak, however, the performance of the exclusion core and its relationship with the strong core is unknown. We examine these questions in this subsection.
We first define the exclusion core for the housing market model. Given an allocation $\mu$, every coalition $C$ directly controls their endowments $\omega(C)$. Moreover, $C$ indirectly controls the endowments of those who occupy $\omega(C)$, and this control right can be extended to more agents. Formally, the set of objects (directly or indirectly) controlled by a coalition $C$ in an allocation $\mu$ is defined to be
@@ 3.2 Balbuzanov and Kotowski’s exclusion core
Definition 4 (balbuzanov2019endowments).
@@
Unlike the strong core, unaffected agents are precluded from joining exclusion blocking coalitions. balbuzanov2019endowments argue that this is necessary to ensure the nonemptiness of the exclusion core in their model. However, unlike the weak core, this restriction does not prevent the exclusion core from achieving Pareto efficiency, since every Pareto inefficient allocation is exclusion blocked by the set of agents who are strictly better off in a Pareto improvement.
Our first finding is that, the exclusion core and the strong core no longer coincide in the housing market model under weak preferences.
@@
In $^{\prime}_{I}$, the strong core = $\emptyset$, while the exclusion core = $\{\mu,\sigma^{\prime}\}$. In either $\mu$ or $\sigma$, because $2$ cannot be made strictly better off, he cannot join any exclusion blocking coalition. Then, $1$ and $3$ cannot evict each other.
Second, we find that the exclusion core can also be empty under weak preferences.
@@
In any individually rational and Pareto efficient allocation, the three agents must trade objects in a cycle, leading to $\mu$ or $\sigma$. Thus, in either $\mu$ or $\sigma$, every agent controls all objects. So, $3$ can exclusion block $\mu$ via $\sigma$, and $1$ can exclusion block $\sigma$ via $\mu$. So, the exclusion core = $\emptyset$.
@@
We provide a necessary and sufficient condition for the exclusion core to be nonempty.
@@
[…]
The partition $T=(T_{1},T_{2},\ldots,T_{t})$ in Proposition 1 does not need to be a TTS. That is, the most preferred objects of $T_{k}\in T$ among $O\backslash\bigcupop\displaylimits_{\ell=1}^{k-1}\omega(T_{\ell})$ do not need to be owned exclusively by $T_{k}$. Therefore, the above condition is more relaxed than the corresponding condition for the strong core in Proposition 3.1. This implies that the exclusion core is more often nonempty than the strong core. Example 1 has shown that when the strong core is empty, the exclusion core may be nonempty.
Proposition 2 is the main result of this section.
@@
(1) The exclusion core is a superset of the strong core and a subset of the weak core. (2) It coincides with the strong core when the strong core is nonempty. (3) However, the strong core and the weak core are equivalence-closed, whereas the exclusion core is not.

### v3

@@ 2 The housing market model
An allocation is a one-to-one mapping $\mu$ from $I$ to $O$ where $\mu(i)$ denotes the object received by $i$. An allocation $\mu$ is individually rational if, $\forall i\in I$, $\mu(i){}_{i}\omega(i)$. Two allocations $\mu$ and $\mu^{\prime}$ are equivalent (in terms of agents’ welfare) if, $\forall i\in I$, $\mu(i)\sim_{i}\mu^{\prime}(i)$. An allocation $\sigma$ is a Pareto improvement over $\mu$ for a coalition $C$ if, $\forall i\in C$, $\sigma(i){}_{i}\mu(i)$, and for some $j\in C$, $\sigma(j)\succ_{j}\mu(j)$. An allocation is Pareto efficient if it does not have a Pareto improvement for all agents.
We define some useful notations. Given any $i\in I$, for every $O^{\prime}\subseteq O$, $B_{i}(O^{\prime})=\{o\in O^{\prime}:o{}_{i}o^{\prime}\text{ for all }o^{\prime}\in O^{\prime}\}$ is the set of $i$’s favorite objects among $O^{\prime}$. For every $o\in O$, $\mathcal{I}_{i}(o)=\{o^{\prime}\in O:o^{\prime}\sim_{i}o\}$ is the set of objects that $i$ views as indifferent with $o$. For every coalition $C$ and every allocation $\mu$, $\mu(C)=\cup_{i\in C}\{\mu(i)\}$ is the set of objects assigned to $C$. By viewing $\omega$ as an initial allocation, $\omega(C)$ is the set of objects initially owned by $C$. For any two allocations $\mu$ and $\sigma$, we call an agent unaffected from $\mu$ to $\sigma$ if he receives indifferent objects in the two allocations. Then, $C_{\sigma\sim\mu}=\{i\in C:\sigma(i)\sim_{i}\mu(i)\}$ is the set of unaffected members of a coalition $C$ from $\mu$ to $\sigma$, and $C_{\sigma>\mu}=\{i\in C:\sigma(i)\succ_{i}\mu(i)\}$ is the set of better-off members of $C$ from $\mu$ to $\sigma$.
Let $\mathcal{M}$ denote the set of markets. In every $M\in\mathcal{M}$, let $\mathcal{A}(M)$ denote the set of allocations and $2^{\mathcal{A}(M)}$ the power set of $\mathcal{A}(M)$. A solution is a correspondence $f:\mathcal{M}\rightarrow\bigcupop\displaylimits_{M\in\mathcal{M}}2^{\mathcal{A}(M)}$ such that, for every $M\in\mathcal{M}$, $f(M)\in 2^{\mathcal{A}(M)}$. We allow $f(M)$ to be empty for some $M$. A solution $f$ is Pareto efficient if, for every $M\in\mathcal{M}$, if $f(M)$ is nonempty, all elements of $f(M)$ are Pareto efficient allocations in $M$.
We introduce the following property as a criterion to assess solution concepts. It requires that a solution is “closed” with respect to the equivalence relation between allocations: if one allocation belongs to the solution, any other equivalent allocation also belongs to the solution. This property is intuitive for coalition blocking, as agents should base their blocking decisions solely on their welfare.
@@
A solution $f$ is equivalence-closed if, in every market $M$, for every two equivalent allocations $\mu$ and $\mu^{\prime}$, $\mu\in f(M)$ if and only if $\mu^{\prime}\in f(M)$.
Shapley and Scarf (1974) introduce the TTC mechanism for the housing market model. Taking strict preferences as inputs, TTC proceeds as follows: in each step, each remaining agent points to his most preferred object and each remaining object points to its owner; then, the agents in each generated cycle exchange their endowments and then leave; this algorithm stops when all agents have left. When agents have strict preferences, the outcome of TTC is unique. When agents have weak preferences, we need to break preference ties before running the algorithm. Depending on different tie-breaking rules, the algorithm may find different allocations for the same preference profile. All those allocations are called the outcomes of TTC under weak preferences.
@@ 2.1 Strong core and weak core
In the standard definition of the core, a coalition blocks an allocation if its members can benefit from a reallocation of their endowments among themselves. Depending on whether all coalition members must strictly benefit, the definition has two variants.
@@
If we replace the first condition by $\forall i\in C$, $\sigma(i)\succ_{i}\mu(i)$,” then we say that $\mu$ is strongly blocked by $C$ via $\sigma$. The weak core consists of allocations that are not strongly blocked.
@@
The strong core is a subset of the weak core. Under strict preferences, the strong core is nonempty and consists of a unique allocation, which is the unique outcome of TTC. Under weak preferences, however, the strong core may be empty, but if nonempty, it must be Pareto efficient. No matter preferences are strict or not, the weak core is nonempty, but it may contain allocations that are not Pareto efficient. This difference arises because unaffected agents are excluded from strong blocking coalitions, yet their participation is necessary for ensuring Pareto efficiency. See the following example for an illustration.
@@
Three agents $1,2,3$ respectively own three objects $a,b,c$. In the following tables, the left one lists agents’ endowments and the allocations under our examination. The other two tables list two preference profiles.
@@
(Strict preferences) In I, since agents’ favorite objects are distinct, $\sigma$ is the unique Pareto efficient allocation. The weak core = $\{\mu,\sigma,\delta\}$, while the strong core = $\{\sigma\}$. $\mu$ and $\delta$ are weakly blocked by $\{1,2,3\}$ via $\sigma$. In contrast, $\mu$ is not strongly blocked because, given that $2$ cannot be made strictly better off and thus cannot join any blocking coalition, $\{1,3\}$ has no rights to reallocate $b$. Similarly, $\delta$ is not strongly blocked because, given that $3$ cannot be made strictly better off and thus cannot join any blocking coalition, $\{1,2\}$ has no rights to reallocate $c$.
(Weak preferences) In $^{\prime}_{I}$, both $1$ and $3$ most prefer $2$’s endowment and then prefer their own endowment, while $2$ is indifferent between all objects. So, $\mu$ and $\delta$ are the individually rational and Pareto efficient allocations for $^{\prime}_{I}$. In $\mu$, $1$ and $2$ exchange endowments, while in $\delta$, $3$ and $2$ exchange endowments.
The weak core = $\{\omega,\mu,\delta\}$. $\omega$ is not Pareto efficient, but it is not strongly blocked because $2$ cannot be made strictly better off and thus cannot join any blocking coalition.
In contrast, the strong core = $\emptyset$. $\mu$ and $\delta$ are the only candidates for the strong core. But $\mu$ is weakly blocked by $\{2,3\}$ via $\delta$, while $\delta$ is weakly blocked by $\{1,2\}$ via $\mu$. Here, because $2$ is indifferent between $a$ and $c$, 22 alternates between $1$ and $3$ in forming blocking coalitions.
Quint and Wako (2004) provide a necessary and sufficient condition for the strong core to be nonempty. Their condition uses the following concepts.
@@
Proposition 0 (Quint and Wako, 2004).
@@
A corollary of this result is that, when the strong core is nonempty, all of its elements are equivalent, and any allocation equivalent to its elements belongs to the strong core. So, the strong core is equivalence-closed. It is easy to prove that the weak core is also equivalence-closed (see Proposition 1).
3 Balbuzanov and Kotowski’s exclusion core
Assuming strict preferences, Balbuzanov and Kotowski (2019) propose the exclusion core in a general model encompassing the housing market model as a special case. The exclusion core and the strong core do not include each other in the general model, but they coincide in the housing market model under strict preferences. These results suggest that we may treat the exclusion core as a new solution for the housing market model. However, under weak preferences, the performance of the exclusion core and its relationship with the strong core is unknown. We examine these questions in this section.
We first apply BK’s definition of the exclusion core to the housing market model. Given an allocation $\mu$, every coalition $C$ directly controls their endowments $\omega(C)$ and indirectly controls the endowments of those who occupy $\omega(C)$, and this indirect control right can be extended to more agents. Formally, in an allocation $\mu$, the set of objects, directly or indirectly, controlled by a coalition $C$ is defined to be
@@ 3 Balbuzanov and Kotowski’s exclusion core
Definition 3 (Balbuzanov and Kotowski, 2019).
@@
Unlike the strong core, unaffected agents are precluded from exclusion blocking coalitions. BK show that this is necessary to ensure the nonemptiness of the exclusion core in their model. However, unlike the weak core, this restriction does not obstruct Pareto efficiency of the exclusion core. Every Pareto inefficient allocation is trivially exclusion blocked by the set of better-off agents in a Pareto improvement, since no agents are worse off and, therefore, the condition (2) of Definition 3 does not apply.
Our first finding is that the exclusion core and the strong core no longer coincide in the housing market model under weak preferences.
@@
In $^{\prime}_{I}$, the strong core = $\emptyset$, while the exclusion core = $\{\mu,\delta\}$. In either $\mu$ or $\delta$, because $2$ cannot be made strictly better off, he cannot join any exclusion blocking coalition. Then, $1$ or $3$ alone cannot evict each other.
Second, the exclusion core can also be empty under weak preferences.
@@
In any individually rational and Pareto efficient allocation, the three agents must trade objects in a cycle, leading to $\mu$ or $\sigma$. So, in either $\mu$ or $\sigma$, every agent controls all objects. Then, $3$ can exclusion block $\mu$ via $\sigma$, and $1$ can exclusion block $\sigma$ via $\mu$. So, the exclusion core = $\emptyset$.
@@
Lemma 1 provides a necessary condition and a different sufficient condition for an allocation to be in the exclusion core. [^6] A single necessary and sufficient condition for a nonempty exclusion core is still an open question.
Lemma 1.
[…]
  1. 1.
    An allocation $\mu$ is in the exclusion core only if $\mu$ is Pareto efficient and there exists a partition of agents $T=(T_{1},T_{2},\ldots,T_{t})$ such that, for every $T_{k}\in T$ and every $i\in T_{k}$, $\mu(T_{k})=\omega(T_{k})$ and $\mu(i)\in B_{i}\big(O\backslash\bigcupop\displaylimits_{\ell=1}^{k-1}\omega(T_{\ell})\big)$.
  2. 2.
    An allocation $\mu$ is in the exclusion core if $\mu$ is Pareto efficient and there exists a partition of agents $T=(T_{1},T_{2},\ldots,T_{t})$ such that, for every $T_{k}\in T$ and every $i\in T_{k}$, $\mu(T_{k})=\omega(T_{k})$ and $\mu(i)\in B_{i}\big(O\backslash\bigcupop\displaylimits_{\ell=1}^{k-1}\omega(T_{\ell})\big)$, and, additionally, for any allocation $\sigma$ such that $I_{\sigma>\mu}\neq\emptyset$, there exists $T_{k}\in T$ such that $T_{k}\cap I_{\mu>\sigma}\neq\emptyset$ and $k<k^{\prime}$ for every $T_{k^{\prime}}\in T$ such that $T_{k^{\prime}}\cap I_{\sigma>\mu}\neq\emptyset$.
In Lemma 1, the partition $T=(T_{1},T_{2},\ldots,T_{t})$ does not need to be a TTS. That is, each $T_{k}$ can allocate their endowments among themselves in a way such that each $i\in T_{k}$ obtains one of his most preferred objects among $O\backslash\bigcupop\displaylimits_{\ell=1}^{k-1}\omega(T_{\ell})$, but $i$’s most preferred objects among $O\backslash\bigcupop\displaylimits_{\ell=1}^{k-1}\omega(T_{\ell})$ do not need to be owned exclusively by $T_{k}$.
The “only if” result in Lemma 1 implies that if the exclusion core is nonempty and $\mu$ is its element, then $\mu$ must be a Pareto efficient outcome of TTC. Specifically, $\mu(T_{k})=\omega(T_{k})$ means that the agents in $T_{k}$ must exchange their endowments in one or several cycles. So, $\mu$ can be produced by TTC via breaking preference ties in a way such that the cycles in the TTC procedure are generated in an order consistent with the order of groups in $(T_{1},T_{2},\ldots,T_{t})$. With this result, to verify whether the exclusion core is nonempty, we can search for its elements by restricting attention to outcomes of TTC. If all outcomes of TTC are not Pareto efficient, or if all Pareto efficient outcomes of TTC are exclusion blocked, it means that the exclusion core is empty.
The “if” result in Lemma 1 allows us to prove that the strong core is a subset of the exclusion core. By Proposition 2.1, if the strong core is nonempty and $\mu$ is its element, there exists a TTS $T^{*}=(T_{1},T_{2},\ldots,T_{t^{*}})$ such that, for every $T_{k}\in T^{*}$ and every $i\in T_{k}$, $\mu(T_{k})=\omega(T_{k})$ and $\mu(i)\in B_{i}(O\backslash\bigcupop\displaylimits_{\ell=1}^{k-1}\omega(T_{\ell}))$. Then, for any allocation $\sigma$ such that $I_{\sigma>\mu}\neq\emptyset$, let $k^{\prime}$ be the smallest $\ell$ such that $T_{\ell}\cap I_{\sigma>\mu}\neq\emptyset$. For every $i\in T_{k^{\prime}}\cap I_{\sigma>\mu}$, it must be that $\sigma(i)\in\bigcupop\displaylimits_{\ell=1}^{k^{\prime}-1}\omega(T_{\ell})$. Then, there must exist $j\in\bigcupop\displaylimits_{\ell=1}^{k^{\prime}-1}T_{\ell}$ such that $\sigma(j)\notin\bigcupop\displaylimits_{\ell=1}^{k^{\prime}-1}\omega(T_{\ell})$. But since $T^{*}$ is a TTS, $j$ must be worse off under $\sigma$. So, there exists $k<k^{\prime}$ such that $j\in T_{k}\cap I_{\mu>\sigma}$, which satisfies the “if” condition in Lemma 1. So, $\mu$ is in the exclusion core.
Proposition 1 presents our main findings for the exclusion core.
@@
(1) The exclusion core is a (possibly empty) subset of the weak core and a superset of the strong core. (2) The strong core and the weak core are equivalence-closed, whereas the exclusion core is not.
Example 1 has shown that when the strong core is empty, the exclusion core may be nonempty. Together with Proposition 1, it implies that the exclusion core is more often nonempty than the strong core. In the next subsection, we prove that the exclusion core is nonempty in a special case of the housing market model in which an agent views any two objects as indifferent if and only if all agents view the two objects as indifferent. This setting can be modeled as a problem in which objects are classified into types and agents view the different copies of each object type as indifferent.
3.1 Housing market with multi-unit objects
This subsection considers a special setting called housing markets with multi-unit objects. In every such market, there exists a finite set of object types, denoted by $\mathcal{O}$, and for each $x\in\mathcal{O}$, there exist a set of copies of $x$, denoted by $O_{x}$. So, $O=\cup_{x\in\mathcal{O}}O_{x}$. Each $i\in I$ owns an object $\omega(i)\in O$ and has a preference relation i over $O$ such that, for every two distinct objects $o$ and $o^{\prime}$, $o\sim_{i}o^{\prime}$ if and only if $\{o,o^{\prime}\}\subseteq O_{x}$ for some $x\in\mathcal{O}$; that is, $o$ and $o^{\prime}$ are different copies of an object type $x$. Let $I_{x}$ denote the set of the owners of the copies of $x$.
In this special setting, the strong core may still be empty. In Example 3, $a$ and $c$ can be viewed as two copies of an object type. The strong core is empty in the example.
However, we prove that the exclusion core is nonempty in this special setting.
@@ 3.1 Housing market with multi-unit objects
In every housing market with multi-unit objects, the exclusion core is nonempty and coincides with the set of TTC outcomes.
To prove this result, we prove that all TTC outcomes in this special setting are Pareto efficient and satisfy the “if” condition in Lemma 1. So, they belong to the exclusion core. On the other hand, the “only if” result of Lemma 1 suggests that all elements of the exclusion core are TTC outcomes. So, the exclusion core equals the set of TTC outcomes.